OpenStudy (anonymous):

Given: x2 - y2 - 4x - 6y - 6 = 0. Write the standard form of the hyperbola. 1) Find the center 2) Find the vertices 3) Find the foci 4) Find the slope of the asymptotes 5) Determine whether the transverse access is vertical or horizontal.

OpenStudy (anonymous):

\[(x ^{2}-4x)-(y ^{2}+6y)=6\] \[(x ^{2}-4x+4-4)-(y ^{2}+6y+9-9)=6\] \[(x-2)^{2}-(y+3)^{2}-4+9=6\] \[(x-2)^{2}-(y+3)^{2}=1\]

OpenStudy (anonymous):

great so the center is (2, -3)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

could you please explain how to find the other stuff? im lost

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

??

OpenStudy (anonymous):

you can convert it to standard form by putting:: x=X+2 and y=Y-3 you get: \[X ^{2}-Y ^{2}=1\]

OpenStudy (anonymous):

how do you find the vertices?

OpenStudy (anonymous):

the coordinates of vertices wrt to new axes are (\[(X=\pm a,0)ie(X=\pm 1,0)\]

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

its ok i understand

OpenStudy (anonymous):

\[(X=\pm a,Y=\pm 0)\] \[x-2=\pm 1,y+3=0\] so the vertices are (3,-3) and (1,-3)

OpenStudy (anonymous):

foci are \[(X=\pm ae,Y=0) \] where X=x-2 and Y=y+3

OpenStudy (anonymous):

e stands for eccentricity

OpenStudy (anonymous):

\[e=\sqrt{1+(b ^{2}/a ^{2})}\]

OpenStudy (anonymous):

got till here????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

asymptotes are:: \[Y=\pm (b/a)X\] where X=x-2 and Y=y+3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

slope: -(coefficient of x)/(coefficient of y)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

equation of transverse axis is y=0

OpenStudy (anonymous):

@kenneyfamily